# Minimax rates of convergence for nonparametric regression under adversarial attacks

### v1

@@ 4.2 Anisotropic Hölder class
In practice, one of the typically desired properties of a regression function is that it is invariant or robust against perturbations of an input in some specific directions. For example, in image classification tasks, the target function may be invariant against the spatial shift of an input image (Simard et al.,, 2003; Krizhevsky et al.,, 2012). In the context of autonomous driving, the parking sign recognition model is robust to certain weather changes, such as snowy days, because snow always accumulates on the upper side of the parking sign.
@@
$\mathcal{F}_{2}(\beta,L)\triangleq\left\{f:[0,1]^{d}\to\mathbb{R}:\left|f(x)-f(z)\right|\leq L_{1}|x_{1}-z_{1}|^{\beta_{1}}+\cdots+L_{d}|x_{d}-z_{d}|^{\beta_{d}}\right\},$ (4.6)
@@ 5 Discussion
In this paper, we focus on the nonparametric regression problem under the adversarial attacks and examine the minimax rates of convergence in the sup-norm. Unlike the minimax analysis for the specific models in Dan et al., (2020) and Xing et al., 2021b , the results established in this paper are of a general nature. They are applicable across diverse regression function classes and arbitrary additive perturbation sets. We show that the minimax rate in the adversarial setting exhibits a modular form, which equals the standard minimax rate in the absence of an adversary, plus the maximum deviation of the true function value within the perturbation set. Applying the general results to specific models is straightforward: it entails determining the standard minimax rate and calculating the largest Lipschitz constant of the functions in the target class. We further investigate two nonparametric function classes, illuminating the impacts of the perturbation structure on the adversarial minimax rates.
It should be pointed out that the proposed plug-in estimation procedure in this paper is nonadaptive, since it depends on information about the unknown perturbation set $\Delta_{n}$. In the context of practical applications, an important direction for future research is to develop estimation procedures that are both adaptive across different function classes and unknown perturbation sets. Another direction is deriving the minimax rates in the general $L_{p}$-norm under the adversarial attacks. In the standard setting, it is well-known that the metric entropy of the regression function class plays a fundamental role in determining the minimax rates of convergence (LeCam,, 1973; Birgé,, 1986; Yatracos,, 1985; Yang and Barron,, 1999). Extending these general theories to the adversarial setting is of great interest.
@@
Appendix A.1 Proofs of the main results
A.1.1 Proof of Lemma 1
Define two sets $\mathcal{S}_{1}=\{(x,x^{\prime}):x\in\mathcal{X},x^{\prime}\in x+\Delta_{n}\}$ and $\mathcal{S}_{2}=\{(x,x^{\prime}):x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X},x^{\prime}\in\mathcal{X}^{\prime}\}$, where $\mathcal{X}^{\prime}=\cup_{x\in\mathcal{X}}(x+\Delta_{n})$. We first prove that $\mathcal{S}_{1}=\mathcal{S}_{2}$. To this end, we just need to show $\mathcal{S}_{1}\subseteq\mathcal{S}_{2}$ and $\mathcal{S}_{2}\subseteq\mathcal{S}_{1}$. We first prove $\mathcal{S}_{1}\subseteq\mathcal{S}_{2}$. For any $(x_{1},x_{1}^{\prime})\in\mathcal{S}_{1}$, we have $x_{1}\in\mathcal{X}$, and there must exist a $\delta_{1}\in\Delta_{n}$ such that $x_{1}^{\prime}=x_{1}+\delta_{1}$. This also means that $x_{1}^{\prime}\in\mathcal{X}^{\prime}$, $x_{1}=x_{1}^{\prime}-\delta_{1}$, and $x_{1}\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}$. Thus we have $(x_{1},x_{1}^{\prime})\in\mathcal{S}_{2}$. From another direction, if $(x_{2},x_{2}^{\prime})\in\mathcal{S}_{2}$, we have $x_{2}\in\mathcal{X}$, and there must exist a $\delta_{2}\in\Delta_{n}$ such that $x_{2}=x_{2}^{\prime}-\delta_{2}$. Thus we get $x_{2}^{\prime}=x_{2}+\delta_{2}\in x_{2}+\Delta_{n}$, and hence $(x_{2},x_{2}^{\prime})\in\mathcal{S}_{1}$. Therefore, $\mathcal{S}_{2}\subseteq\mathcal{S}_{1}$.
@@ A.1.1 Proof of Lemma 1
$\sup_{x\in\mathcal{X}}\sup_{x^{\prime}\in x+\Delta_{n}}\left|f(x)-\widehat{f}(x^{\prime})\right|=\sup_{(x,x^{\prime})\in\mathcal{S}_{1}}\left|f(x)-\widehat{f}(x^{\prime})\right|$ (A.1.1)
and
$\sup_{x^{\prime}\in\mathcal{X}^{\prime}}\sup_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}\left|f(x)-\widehat{f}(x^{\prime})\right|=\sup_{(x,x^{\prime})\in\mathcal{S}_{2}}\left|f(x)-\widehat{f}(x^{\prime})\right|$ (A.1.2)
Assume $\sup_{x\in\mathcal{X}}\sup_{x^{\prime}\in x+\Delta_{n}}|f(x)-\widehat{f}(x^{\prime})|<\sup_{(x,x^{\prime})\in\mathcal{S}_{1}}|f(x)-\widehat{f}(x^{\prime})|$. Then there must exist a pair $(x_{3},x_{3}^{\prime})\in\mathcal{S}_{1}$ such that $\sup_{x\in\mathcal{X}}\sup_{x^{\prime}\in x+\Delta_{n}}|f(x)-\widehat{f}(x^{\prime})|<|f(x_{3})-\widehat{f}(x_{3}^{\prime})|$. On the other hand, based on the definition of $\mathcal{S}_{1}$, we have $x_{3}^{\prime}\in x_{3}+\Delta_{n}$. Thus,
$\sup_{x\in\mathcal{X}}\sup_{x^{\prime}\in x+\Delta_{n}}|f(x)-\widehat{f}(x^{\prime})|\geq\sup_{x^{\prime}\in x_{3}+\Delta_{n}}|f(x_{3})-\widehat{f}(x^{\prime})|\geq|f(x_{3})-\widehat{f}(x_{3}^{\prime})|,$
which is a contradiction. Likewise, we can prove that $\sup_{x\in\mathcal{X}}\sup_{x^{\prime}\in x+\Delta_{n}}|f(x)-\widehat{f}(x^{\prime})|>\sup_{(x,x^{\prime})\in\mathcal{S}_{1}}|f(x)-\widehat{f}(x^{\prime})|$ is also impossible. Therefor, (A.1.1) is proved. The equality (A.1.2) can be proved in a similar manner. We thus skip the similar materials here.
A.1.2 Proof of Theorem 1
Based on the results in Lemma  1, we have
$\begin{split}L_{\Delta_{n}}(f,f^{\prime})&=\sup_{x^{\prime}\in\mathcal{X}^{\prime}}\sup_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}\left|f(x)-f^{\prime}(x^{\prime})\right|\\ &=\sup_{x^{\prime}\in\mathcal{X}^{\prime}}\left[\left|\frac{\sup_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}f(x)+\inf_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}f(x)}{2}-f^{\prime}(x^{\prime})\right|\right.\\ &\left.+\frac{\sup_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}f(x)-\inf_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}f(x)}{2}\right].\end{split}$ (A.1.3)
Thus, the minimum of $L_{\Delta_{n}}(f,f^{\prime})$ is obtained when
@@ A.1.2 Proof of Theorem 1
for any $x^{\prime}\in\mathcal{X}^{\prime}$. And the ideal adversarial risk is given by
$\frac{1}{2}\sup_{x^{\prime}\in\mathcal{X}^{\prime}}\left[\sup_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}f(x)-\inf_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}f(x)\right],$
@@ A.1.3 Proof of Theorem 2
$\begin{split}R_{\Delta_{n}}(f,\widehat{f}_{\mathrm{PI}})&=\mathbb{E}\sup_{x^{\prime}\in\mathcal{X}^{\prime}}\left[\left|\frac{\sup_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}f(x)+\inf_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}f(x)}{2}-\widehat{f}_{\mathrm{PI}}(x^{\prime})\right|\right.\\ &\left.+\frac{\sup_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}f(x)-\inf_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}f(x)}{2}\right].\end{split}$ (A.1.4)
@@
$\begin{split}&\left|\frac{\sup_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}f(x)+\inf_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}f(x)}{2}-\widehat{f}_{\mathrm{PI}}(x^{\prime})\right|\\ &=\left|\frac{\sup_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}f(x)+\inf_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}f(x)}{2}-\frac{\sup_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}\widetilde{f}(x)+\inf_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}\widetilde{f}(x)}{2}\right|\\ &\leq\frac{1}{2}\left|\sup_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}f(x)-\sup_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}\widetilde{f}(x)\right|+\frac{1}{2}\left|\inf_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}f(x)-\inf_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}\widetilde{f}(x)\right|\\ &\leq\sup_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}\left|f(x)-\widetilde{f}(x)\right|+\sup_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}f(x)-\inf_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}f(x).\end{split}$ (A.1.5)
@@
$\begin{split}R_{\Delta_{n}}(f,\widehat{f}_{\mathrm{PI}})&\leq\mathbb{E}\sup_{x^{\prime}\in\mathcal{X}^{\prime}}\sup_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}\left|f(x)-\widetilde{f}(x)\right|+3L_{\Delta_{n}}^{*}(f)\\ &\leq R(f,\widetilde{f})+3L_{\Delta_{n}}^{*}(f),\end{split}$
which completes the proof.
The second part of this theorem is proved by taking upper bound on both sides of (3.5) with respect to $f\in\mathcal{F}_{1}(\beta,L)$ and then using the condition (3.6).
A.1.4 Proof of Theorem 3

### v2

@@ 4.2 Anisotropic Hölder class
In practice, one of the typically desired properties of a regression function or its estimator is that it is invariant or robust against changes or perturbations of an input in some specific directions. For example, in image classification tasks, the target function should be invariant against a spatial shift or rotation of an input image (Simard et al.,, 2003; Krizhevsky et al.,, 2012). In the same spirit, in the context of autonomous driving, a traffic sign recognition model should be trained to be robust to natural variations in severe weather conditions.
@@
$\begin{split}\mathcal{F}_{2}(\beta,L)\triangleq\left\{f:[0,1]^{d}\to\mathbb{R}:\right.&\left|f(x)-f(z)\right|\\ &\left.\leq L_{1}|x_{1}-z_{1}|^{\beta_{1}}+\cdots+L_{d}|x_{d}-z_{d}|^{\beta_{d}}\right\},\end{split}$ (4.6)
@@
5 Simulation studies
In this section, we present several numerical experiments to illustrate the theoretical results established in Sections 34. The data are generated from the model (2.1), where $\mathcal{X}=[0,1]^{2}$, $X$ follows a uniform distribution on $[0,1]^{2}$, and $\xi$ is independent of $X$ and distributed as $N(0,\sigma^{2})$. We consider several regression functions and attack scenarios:
Case 1
$f(x_{1},x_{2})=\sqrt{x_{1}x_{2}}$ with perturbation set $\Delta_{n}=B_{\infty}^{r}$.
Case 2
$f(x_{1},x_{2})=\sqrt{(x_{1}-0.5)^{2}+(x_{2}-0.5)^{2}}$, with $\Delta_{n}=B_{\infty}^{r}$.
Case 3
$f(x_{1},x_{2})=\sqrt{x_{1}}+0.1x_{2}-0.5$, with $\Delta_{n}=[-4r,+4r]\times[-r/4,+r/4]$.
Case 4
$f(x_{1},x_{2})=\sqrt{x_{1}}+0.1x_{2}-0.5$, with $\Delta_{n}=[-r/4,+r/4]\times[-4r,+4r]$.
In each case, $\sigma^{2}$ is adjusted so that the signal-to-noise ratio equals $5$. The attack magnitude $r$ increases from 0 to 0.1. Cases 1–2 serve as two representative examples of isotropic Hölder classes, where the perturbation set is chosen as the $\ell_{\infty}$-ball. In contrast, Cases 3–4 consider regression functions with different degrees of variation along different axes, where the attack magnitudes are also anisotropic.
We consider three competing methods. The baseline method (LP) is the classical local polynomial regression studied in Stone, (1982), Bertin, 2004b , and Gaïffas, (2007) based on the rectangular kernel. We employ a polynomial of degree $\ell=1$ (i.e., local linear regression). In Cases 1 and 2, the bandwidth is set as $h=n^{-1/(0.5+2)}$ and $h=n^{-1/(1+2)}$, respectively. In Cases 3 and 4, we use different bandwidths for different coordinates, setting $h_{1}=n^{-1/(0.5+2)}$ and $h_{2}=n^{-1/(1+2)}$. These choices are theoretically proven to achieve the standard minimax rates in the respective cases.
Building on the LP method, we consider two additional competing methods. The first (PI) follows (3.4), where $\widetilde{f}$ is the LP estimator. The second method is a ridge-type local polynomial estimator (RG), which follows the LP approach but incorporates a ridge penalty with parameter $r^{2}$ on the linear coefficients during the estimation of the LP coefficients. The ridge-type strategy can be seen as an approximation of adversarial training (Ribeiro and Schön,, 2023) and has also been proven to possess desirable robustness properties under several specific setups (Zhang et al.,, 2019; Xing et al., 2021b, ). Figures 12 present the adversarial risk for the three competing methods over 100 simulation replications. In each replication, the adversarial loss is evaluated at 100 uniformly sampled points in $[0,1]^{2}$.
(a) Case 1
(b) Case 2
Figure 1: Adversarial risk for the three competing methods as the attack magnitude increases: panel (a) corresponds to Case 1, and panel (b) corresponds to Case 2.
(a) Case 3
(b) Case 4
Figure 2: Adversarial risk for the three competing methods as the attack magnitude increases: panel (a) corresponds to Case 3, and panel (b) corresponds to Case 4.
From Figures 12, we observe a significant advantage of the PI method over the classical LP method and its ridge-type variant. For instance, in Case 2 with $n=200$ and $r=0.5$, the adversarial risk and its standard error for LP, RG, and PI are 2.73e-2 (0.001), 2.71e-2 (0.001), and 2.21e-2 (0.001), respectively. These results demonstrate that the adversarial plug-in procedure 3.4 achieves a substantial improvement in robustness compared to the other two methods. The patterns depicted in Figures 12 further corroborate the insights discussed in Sections 4.14.2. For example, in Case 1, where the regression function belongs to $\mathcal{F}_{1}(\beta,L)$ with $\beta=1/2$, the adversarial risk curve exhibits a concave shape, consistent with the $r^{1/2}$. In Case 2, the adversarial risk curve is approximately linear as $r$ increases, which aligns with Example 1 that the adversarial loss in this case is dominated by $r$ when $r$ is large. Additionally, Figure 2 reveals that strong attacks along directions with higher variability can significantly degrade the performance of competing methods, supporting the theoretical results presented in Example 2.
Furthermore, although existing literature suggests that ridge-type regularization can enhance adversarial robustness under various modeling frameworks (see, e.g., Zhang et al.,, 2019; Xing et al., 2021b, ), its effectiveness in the context of local nonparametric estimation remains limited. This limitation arises because ridge regularization in RG primarily controls the variation of the LP estimator at a given local point but does not regulate the variation of the estimator across different local points. Consequently, the RG method may still be vulnerable to adversarial attacks under our context.
@@ 6 Discussion
In this paper, we focus on the nonparametric regression problem under the adversarial attacks and examine the minimax rates of convergence in the adversarial sup-norm. Unlike the minimax analysis for the specific models in Dan et al., (2020) and Xing et al., 2021b , the results established in this paper are of a general nature. They are applicable across diverse regression function classes and arbitrary additive perturbation sets. We show that the minimax rate in the adversarial setting exhibits a modular form, which equals the standard minimax rate in the absence of an adversary, plus the maximum deviation of the true function value within the perturbation set. Applying the general results to specific models is straightforward: it entails determining the standard minimax rate and calculating the largest Lipschitz constant of the functions in the target class. We further investigate two nonparametric function classes, illuminating the impacts of the different perturbation sets on the adversarial minimax rates.
It should be pointed out that the proposed adversarial plug-in estimation procedure in this paper is nonadaptive, since it depends on information about the unknown perturbation set $\Delta_{n}$. In the context of practical applications, an important direction for future research is to develop estimation procedures that are both adaptive across different function classes and unknown perturbation sets. Another direction is deriving the minimax rates in the general $L_{p}$-norm under the adversarial attacks. In the standard setting, it is well-known that the metric entropy of the regression function class plays a fundamental role in determining the minimax rates of convergence (LeCam,, 1973; Birgé,, 1986; Yatracos,, 1985; Yang and Barron,, 1999). Extending these general theories to the adversarial setting is of great interest.
@@
Appendix A.1 Proof of Lemma 1
Recall the definition of the adversarial sup-norm loss in (2.2). By the change of variable $x^{\prime}=x+\delta$, we have $x^{\prime}\in(x+\Delta_{n})\cap\mathcal{X}$ since $\delta\in\Delta_{n}$. Therefore, the adversarial loss can be expressed as
$L_{\Delta_{n}}(f,\widehat{f})=\sup_{x\in\mathcal{X}}\sup_{x^{\prime}\in(x+\Delta_{n})\cap\mathcal{X}}\left|f(x)-\widehat{f}(x^{\prime})\right|.$ (A.1.1)
@@ Appendix A.1 Proof of Lemma 1
$\sup_{x\in\mathcal{X}}\sup_{x^{\prime}\in(x+\Delta_{n})\cap\mathcal{X}}\left|f(x)-\widehat{f}(x^{\prime})\right|=\sup_{x^{\prime}\in\mathcal{X}}\sup_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}\left|f(x)-\widehat{f}(x^{\prime})\right|.$ (A.1.2)
Assume that
$\sup_{x\in\mathcal{X}}\sup_{x^{\prime}\in(x+\Delta_{n})\cap\mathcal{X}}|f(x)-\widehat{f}(x^{\prime})|>\sup_{x^{\prime}\in\mathcal{X}}\sup_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}|f(x)-\widehat{f}(x^{\prime})|.$
Then there must exist $x_{1}\in\mathcal{X}$ and $x_{1}^{\prime}\in(x_{1}+\Delta_{n})\cap\mathcal{X}$ such that $|f(x_{1})-\widehat{f}(x_{1}^{\prime})|>\sup_{x^{\prime}\in\mathcal{X}}\sup_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}|f(x)-\widehat{f}(x^{\prime})|$. On the other hand, based on the definition of $(x_{1},x_{1}^{\prime})$, we have $x_{1}^{\prime}\in\mathcal{X}$ and $x_{1}=x_{1}^{\prime}-\delta_{1}$ for some $\delta_{1}\in\Delta_{n}$, implying that $x_{1}\in(x_{1}^{\prime}-\Delta_{n})\cap\mathcal{X}$. This leads to
$|f(x_{1})-\widehat{f}(x_{1}^{\prime})|\leq\sup_{x^{\prime}\in\mathcal{X}}\sup_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}|f(x)-\widehat{f}(x^{\prime})|,$
which is a contradiction. Likewise, we can prove that $\sup_{x\in\mathcal{X}}\sup_{x^{\prime}\in(x+\Delta_{n})\cap\mathcal{X}}|f(x)-\widehat{f}(x^{\prime})|<\sup_{x^{\prime}\in\mathcal{X}}\sup_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}|f(x)-\widehat{f}(x^{\prime})|$ is also impossible. Therefore, (A.1.2) is proved.
Appendix A.2 Proof of Theorem 1
Based on the results in Lemma 1, we have
$\begin{split}L_{\Delta_{n}}(f,f^{\prime})&=\sup_{x^{\prime}\in\mathcal{X}}\sup_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}\left|f(x)-f^{\prime}(x^{\prime})\right|.\\ \end{split}$ (A.2.1)
For any given $x^{\prime}\in\mathcal{X}$, note that
$\begin{split}&\sup_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}\left|f(x)-f^{\prime}(x^{\prime})\right|\\ &=\max\left\{\left|\sup_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}f(x)-f^{\prime}(x^{\prime})\right|,\left|\inf_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}f(x)-f^{\prime}(x^{\prime})\right|\right\}\\ &=\left[\left|\frac{\sup_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}f(x)+\inf_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}f(x)}{2}-f^{\prime}(x^{\prime})\right|\right.\\ &\left.\quad+\frac{\sup_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}f(x)-\inf_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}f(x)}{2}\right],\end{split}$ (A.2.2)
where the first equality follows from the fact that $|f(x)-f^{\prime}(x^{\prime})|$, as a piecewise linear function of $f(x)$, achieves the supremum when $f(x)$ attains either its supremum $\sup_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}f(x)$ or its infimum $\inf_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}f(x)$, and the second equality is established by analyzing the relative values of $\sup_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}f(x)$, $\inf_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}f(x)$, and $f^{\prime}(x^{\prime})$.
Combining (A.2.1) with (A.2.2), we obtain
$\begin{split}L_{\Delta_{n}}(f,f^{\prime})&=\sup_{x^{\prime}\in\mathcal{X}}\left[\left|\frac{\sup_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}f(x)+\inf_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}f(x)}{2}-f^{\prime}(x^{\prime})\right|\right.\\ &\left.\quad+\frac{\sup_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}f(x)-\inf_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}f(x)}{2}\right].\end{split}$ (A.2.3)
Since $f^{\prime}$ appears only in the absolute value term in (A.2.3), the infimum $\inf_{f^{\prime}}L_{\Delta_{n}}(f,f^{\prime})$ is therefore obtained when
@@ Appendix A.2 Proof of Theorem 1
for any $x^{\prime}\in\mathcal{X}$. And the ideal adversarial risk is given by
$\frac{1}{2}\sup_{x^{\prime}\in\mathcal{X}}\left[\sup_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}f(x)-\inf_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}f(x)\right],$
@@ Appendix A.3 Proof of Theorem 2
$\begin{split}R_{\Delta_{n}}(f,\widehat{f}_{\mathrm{PI}})&=\mathbb{E}\sup_{x^{\prime}\in\mathcal{X}}\left[\left|\frac{\sup_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}f(x)+\inf_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}f(x)}{2}-\widehat{f}_{\mathrm{PI}}(x^{\prime})\right|\right.\\ &\left.+\frac{\sup_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}f(x)-\inf_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}f(x)}{2}\right].\end{split}$ (A.3.1)
@@
$\begin{split}&\left|\frac{\sup_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}f(x)+\inf_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}f(x)}{2}-\widehat{f}_{\mathrm{PI}}(x^{\prime})\right|\\ &=\left|\frac{\sup_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}f(x)+\inf_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}f(x)}{2}\right.\\ &\quad\left.-\frac{\sup_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}\widetilde{f}(x)+\inf_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}\widetilde{f}(x)}{2}\right|\\ &\leq\frac{1}{2}\left|\sup_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}f(x)-\sup_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}\widetilde{f}(x)\right|\\ &\quad+\frac{1}{2}\left|\inf_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}f(x)-\inf_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}\widetilde{f}(x)\right|\\ &\leq\sup_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}\left|f(x)-\widetilde{f}(x)\right|.\end{split}$ (A.3.2)
@@
$\begin{split}R_{\Delta_{n}}(f,\widehat{f}_{\mathrm{PI}})&\leq\mathbb{E}\sup_{x^{\prime}\in\mathcal{X}}\sup_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}\left|f(x)-\widetilde{f}(x)\right|+L_{\Delta_{n}}^{*}(f)\\ &\leq R(f,\widetilde{f})+L_{\Delta_{n}}^{*}(f),\end{split}$
where the first inequality follows from (A.3.2)–(A.3.1) and the definition of $L_{\Delta_{n}}^{*}(f)$ in (3.2), and the second inequality follows from
$\sup_{x^{\prime}\in\mathcal{X}}\sup_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}\left|f(x)-\widetilde{f}(x)\right|\leq\sup_{x^{\prime}\in\mathcal{X}}\sup_{x\in\mathcal{X}}\left|f(x)-\widetilde{f}(x)\right|=\sup_{x\in\mathcal{X}}\left|f(x)-\widetilde{f}(x)\right|.$
Thus, we complete the proof of (3.5).
The second part of this theorem is proved by taking upper bound on both sides of (3.5) with respect to $f\in\mathcal{F}$ and then using the condition (3.6). Specifically, we have
$\begin{split}\sup_{f\in\mathcal{F}}R_{\Delta_{n}}(f,\widehat{f}_{\mathrm{PI}})&\lesssim\sup_{f\in\mathcal{F}}R(f,\widetilde{f})+\sup_{f\in\mathcal{F}}L_{\Delta_{n}}^{*}(f)\\ &\asymp\inf_{\widehat{f}}\sup_{f\in\mathcal{F}}R(f,\widehat{f})+\sup_{f\in\mathcal{F}}L_{\Delta_{n}}^{*}(f),\end{split}$
which leads to (3.7).
Appendix A.4 Proof of Theorem 3